OpenStudy (iamstark):

Find out the wavelength of the light emitted if electron jumps from 5th to 2nd orbit in terms of Rydberg constant

OpenStudy (iamstark):

@KL-RC

OpenStudy (iamstark):

@abb0t

OpenStudy (iamstark):

its okay @KL-RC , can u tag someone who might know this

Parth (parthkohli):

\[\dfrac{1}{\lambda} = R\left(\dfrac{1}{n_f^2} - \dfrac{1}{n_i^2}\right)\]

OpenStudy (iamstark):

Thnx @ParthKohli ...then ithink it should be 5/R

Parth (parthkohli):

\[\dfrac{1}{\lambda} = R\left(\dfrac{1}{4} - \dfrac{1}{25}\right)\]\[= R\cdot \dfrac{21}{100}\]Take the reciprocal of both sides...

OpenStudy (iamstark):

yes..approx 5/R

Parth (parthkohli):

OK, yes. I thought you calculated that in exact values and got a little worried. lol

OpenStudy (iamstark):

:)